OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x: f(x) =x-7/x+3 and g(x) = -3x-7/x-1

OpenStudy (anonymous):

f(g(x)) = x let's do this one first. since g(x) took the place of x in f(x) you have to put in g(x) in place of x in equation f(x). What I mean here is... f(g(x))=g(x)-7/g(x)+3=(-3x+7/x+1)-7/(-3x+7/x+1)+3 simplify that equation and you should get f(g(x))=x for the second part g(f(x))=x, do the same thing but now put in f(x) in g(x).

OpenStudy (anonymous):

you have it wrong it is g(x)=-3x-7/x+1 not -3x+7/x+1

OpenStudy (anonymous):

also i had it wrong suppose to be x-1 under the -3x-7

OpenStudy (jhannybean):

\[f(x)= \cfrac{x-7}{x+3}\]\[g(x)=\cfrac{-3x-7}{x+1}\]\[f(g(x)) = \cfrac{\cfrac{-3x-7}{x+1} -7}{\cfrac{-3x-7}{x+1} +1}\]\[ f(g(x)) =\cfrac{\cfrac{-3x-7-7x-7}{x+1}}{\cfrac{-3x-7+x+1}{x+1}}\]\[f(g(x)) = \cfrac{\cfrac{-10x-14}{x+1}}{\cfrac{-2x-6}{x+1}} \]\[f(g(x)) = \cfrac{-10x-14}{-2x-6}\]\[ f(g(x)) = \cfrac{-2(5x+7)}{-2(x+3)}\]\[f(g(x)) = \cfrac{5x+7}{x+3}\]

OpenStudy (jhannybean):

omg i messed up....

OpenStudy (anonymous):

try it again?

OpenStudy (anonymous):

made a typo there but I used the same idea as Jhannybean

OpenStudy (jhannybean):

\[f(g(x)) = \cfrac{\cfrac{-3x-7}{x+1} -7}{\cfrac{-3x-7}{x+1} +3}\]\[f(g(x)) =\cfrac{\cfrac{-3x-7-7x-7}{x+1}}{\cfrac{-3x-7+3x+3}{x+1}}\]\[f(g(x))= \cfrac{-10x-14}{-4}\]\[f(g(x))= \cfrac{5x-7}{2}\] and you will get the same result solving \(g(f(x))\)

OpenStudy (anonymous):

I also have a problem with that one because it is as shown \[g(f(x))=\frac{ -3(\frac{ x-7 }{ x+3 })-7 }{ \frac{ x-7 }{ x+3 }-1 }\]

OpenStudy (anonymous):

And i got no idea how to reduce that

OpenStudy (raden):

keep spirit :) @Jhannybean

OpenStudy (jhannybean):

Maybe theyre not inverses of each other.... hmm... lets see!

OpenStudy (jhannybean):

\[g(f(x)) = \cfrac{-3\left(\cfrac{x-7}{x+3}\right)-7}{\cfrac{x-7}{x+3}+1}\]\[g(f(x)) = \cfrac{\cfrac{-3x+21-7x-21}{x+3}}{\cfrac{x-7+x+3}{x+3}}\]\[g(f(x)) = \cfrac{-3x-7x}{2x-4}\]\[g(f(x)) = \cfrac{-10x}{2x-4}\]\[g(f(x)) = \cfrac{-10x}{2(x-2)}\]\[g(f(x))= \cfrac{-5x}{x-2}\]

OpenStudy (anonymous):

so are they inverses?

OpenStudy (jhannybean):

I don't think they are.....

OpenStudy (anonymous):

hmm unusual. Most the questions i was askedi n the past involving this they were

OpenStudy (jhannybean):

I am kind of unsure atm xD

OpenStudy (anonymous):

Eh if all else fails i can submit my assignment and just get a 34/35 asuming i got rest all ok

OpenStudy (anonymous):

ok well i put this down and see wht i get

OpenStudy (anonymous):

thanks!

OpenStudy (jhannybean):

One minute...

OpenStudy (anonymous):

ok I will wait

OpenStudy (jhannybean):

http://fooplot.com/plot/fwj9gaersw they don't look like inverses of eachother, perhaps @RadEn can help clarify this?

OpenStudy (anonymous):

no they dont...this is quite odd to be honest xD

OpenStudy (raden):

yes, f and g are inverses. maybe just wrong in simplication :)

OpenStudy (anonymous):

well when simplified doing f(g(x)) and g(f(x)) they should both equal x

OpenStudy (raden):

yes, i got that

OpenStudy (anonymous):

but it looks too complicated to even think that they do xD that is the problem and jhannybean's work looks all ok

OpenStudy (jhannybean):

Can you point out where i went wrong? xD

OpenStudy (raden):

g(f(x)) = ...../(x-7)/(x+3) + 1 the last should -1 not + 1

OpenStudy (anonymous):

ahh yes

OpenStudy (anonymous):

hopefully i will see xD

OpenStudy (anonymous):

Ok got first one to equal x now all i got is the g(f(x))

OpenStudy (raden):